A capacitor circuit has been charged up to 12V and the instantaneous voltage is ๐’— = ๐Ÿ๐Ÿ (๐Ÿ โˆ’ ๐’† โˆ’ ๐’• ๐œ)

The tasks are to:ย 

a) Draw a graph of voltage against time for ๐‘ฃ = 12๐‘‰ and ๐œ = 2๐‘ , between ๐‘ก = 0๐‘  and ๐‘ก = 10๐‘ .

b) Calculate the gradient at ๐‘ก = 2๐‘  and ๐‘ก = 4๐‘ 

c) Differentiate ๐’— = ๐Ÿ๐Ÿ (๐Ÿ โˆ’ ๐’† โˆ’ ๐’• ๐œ) and calculate the value of ๐‘‘๐‘ฃ ๐‘‘๐‘ก at ๐‘ก = 2๐‘  and ๐‘ก = 4๐‘ .

d) Compare your answers for part b and part c.ย 

e) Calculate the second derivative of the instantaneous voltage ( ๐‘‘^2๐‘ฃ/๐‘‘๐‘ก^2 ).ย 


1
Expert's answer
2022-02-22T13:52:21-0500

Given: The equation for instantaneous voltage is

"V=12(1-e^{-t/\tau})"

a)


"V=12(1-e^{-t/2}), t\in [0, 10]"


b)

"grad|_{t=2s}\approx\dfrac{12(1-e^{-(2+\Delta t/2})-12(1-e^{-2/2})}{\Delta t}"

"\approx\dfrac{12e^{-1}(1-e^{-\Delta t/2})}{\Delta t}\approx6/e\approx2.2073\ V/s"

"grad|_{t=4s}\approx\dfrac{12(1-e^{-(4+\Delta t/2})-12(1-e^{-4/2})}{\Delta t}"

"\approx\dfrac{12e^{-2}(1-e^{-\Delta t/2})}{\Delta t}\approx6/e^2\approx0.8120\ V/s"

c)ย 

"V'(t)=12(-1\/2)(-e^{-t\/2})=6e^{-t\/2}"


ย 

"V'(2)=6/e\approx2.2073\ V/s"

"V'(4)=6/e^2\approx0.8120\ V/s"

d) The result is the same.


e)


"V''(t)=-3e^{-t\/2}"